Author Archives: Isaac Bakerman

140827 LML Observations

Isaac arrived at the collector at 9:30am. The doors to the collector had been deployed however, no sample had been collected. The collector was then disassembled for cleaning. This was preformed by utilizing 409 and DI water. Upon completion of this process the collector was then reassembled and reset for the next fog event.

140830 LML Observations

Isaac arrived at LML at 10:20 am. The doors to the collector had been deployed and a sample had been collected in the sample jar. The sample measured approximately 50 ml. The sample was relatively clear with a small amount of sediment at the bottom. The sample was then removed from the collector and placed in a small labeled bag. The collector was then dissembled and cleaned with 409 and DI water. After the collector was reassembled a blank was then taken which measured approximately 200ml. The collector was then reset for the next fog event.

140820 UCSC Observations

Isaac arrived at ISB at 9:45am. Upon arrival there was a light drizzle coming down. The doors to the collector had been deployed and the big fag was on. However, upon examination of the collection jar no sample was present. The collector was then disassembled and its components as well as the collector body were washed with 409 and DI water. The system was then reset for the next fog collection event. In addition due to the presence of precipitation a passive rain collector was set up in order to collect any rainfall.

140817 LML Observations

Isaac arrived at LML at 9:30am. The doors to the collector had been deployed but no sample was present in the collection jar. The collector was then disassembled and the body of the collector as well as its components were sprayed initially with 409 followed by a rinse with DI water. The collector was then reassembled for the next fog event.

140816 LML Observations

Isaac arrived at LML at 10:30 am. The doors to the collector had been deployed but no sample was present in the collection jar. The collector was then disassembled and washed with 409 and DI water. Upon completion of this process the collector was reassembled for the next fog event.

LML Observations 140814

Isaac arrived at LML at 9:30 am. The doors to the collector had been deployed and there was a sample present in the collection jar. the sample was approximately 25ml and had a light pale yellow coloration. the sample was removed and placed in a labeled bag. the collector was then disassembled and its components were washed with 409 followed by DI water. After the collector had been cleaned a blank was taken. the blank sample measured a volume of approximately 200ml. After the blank was taken the system was reset for the next fog event.

140813 LML Observations

Isaac arrived at LML at 9:50 am. The doors to the collector had been deployed but no sample was present. The collector was then disassembled and the collector components were cleaned with 409 and DI water. The collector body was then cleaned with 409 and rinsed the DI water as well. The collector was then reassembled and reset for the next fog event.

140811 LML Observations

Isaac arrived at the LML site at 9:40am. The doors to the collector had been deployed and there was a small sample present but not enough for further analysis. The small sample was discarded. The collector was then disassembled and washed with 409 followed by DI water. A significant amount of bugs had accumulated on the strand frames which were removed during the washing process. The body of the collector was then washed and rinsed with DI. The collector was then reset for the next fog event.

140808 UCSC Observations

Isaac arrived at ISB at 9:47am. The doors to the collector had opened and there was a sample in the collection jar. The sample appeared clear with a small amount of sediment in the bottom of the jar. The sample was then removed and placed in a labeled bag. The collector was then disassembled and its components were cleaned with 409 and DI water. The collector body was then cleaned with 409 and DI water as well. The collector was then reassembled with the cleaned components. Upon reassembly a blank was taken by spraying the string frames and allowing the DI water to drain into a newly placed jar. This blank was then removed and placed in a labeled bag. The system was then reset for the next fog event.

140804 LML Observations

Isaac arrived at LML at 9:30am. The doors to the collector had deployed and a tiny sample of approximately 5ml was visible in the collection jar. Since the sample was of minimal volume it was discarded. The collector was then sprayed with 409 and rinsed with DI water. the collector was additionally disassembled and its components cleaned with 409 and DI water. The collector was then reassembled for the next fog event.